To buy both a new car and new house Tina sought two loans totaling $260,394
The simple interest rate on the first loan was 3.7percent
The interest rate on the second loan was 8.8 percent. At the end of the first year
Tina paid a combined interest payment of $22,611.73. what was the amounts of the two loans?

Answer by jorel1380(3719)

m+n=260394
.037m+.088n=22611.73
37m+88n=22611730
37m+37n=9634578
51n=12977152
n=25443.96
m=234950.04
Tina's first loan was for $234950.04. Her second loan was for $25443.96..